# Trade-offs in Data Memorization via Strong Data Processing Inequalities

### v1

@@ D.3 Sparse Boolean memorization upper bound
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{sB}})=I(h,X_{1:n}\mid\theta)\overset{(h\perp\!\!\!\perp\theta\,|\,X_{1:n})}{\leq}I(h,X_{1:n})\leq H(h)=\widetilde{O}\left(\frac{d}{2^{2n}}\right)~{}.$
Let $\hat{S}\subset[d]$ be the subset that includes coordinates $j\in[d]$ in which the dataset is constant, namely for which $(X_{i})_{j}=(X_{i^{\prime}})_{j}$ is the same for all $i,i^{\prime}\in[n]$. Furthermore, let $\hat{S}^{\ell}$ be the first $\ell$ coordinates in $\hat{S}$, and define
@@
$\ell=\Theta\left(\frac{d}{2^{2n}}\right),~{}~{}~{}t=\sqrt{2\log(200)\ell}=\Theta(\sqrt{\ell}),$
this classifier satisfies the desired properties. We note that the memorization upper bound readily follows by the description length argument above, since all that needs to be stored are the $\ell$ bits of $(X_{1})_{\hat{S}^{\ell}}$ alongside their corresponding indices in $[d]$, which requires $O(\ell\log(d))$ memory.
It remains to show that $h$ is a high accuracy classifier. We first note that in the null case $\mathcal{P}_{0}=\mathcal{U}(\{\pm 1\}^{d})$, so it holds that
$\mathop{\mathbb{E}}_{X\sim\mathcal{P}_{0}}\langle X_{\hat{S}^{\ell}},(X_{1})_{\hat{S}^{\ell}}\rangle=0~{},$
@@
$\Pr\left[\langle X_{\hat{S}^{\ell}},(X_{1})_{\hat{S}^{\ell}}\rangle\geq t\right]=\Pr\left[\langle X_{\hat{S}^{\ell}},(X_{1})_{\hat{S}^{\ell}}\rangle\geq\sqrt{2\log(200)\ell}\right]\leq 0.005~{}.$ (10)
@@
$\mathop{\mathbb{E}}_{X\sim\mathcal{P}_{\theta}}\langle X_{\hat{S}^{\ell}},(X_{1})_{\hat{S}^{\ell}}\rangle=|\hat{S}^{\ell}\cap S|~{}.$ (11)
@@
Moreover, by symmetry, choosing the first $\ell$ coordinates in $\hat{S}$ into $\hat{S}^{\ell}$ is equivalent to sub-sampling $\ell$ coordinates among the constant coordinates, and therefore
$\mathop{\mathbb{E}}|\hat{S}^{\ell}\cap S|=\ell\cdot\Pr[j\in S\mid j\in\hat{S}]=\ell\cdot\left(\frac{\nu}{\nu+(1-\nu)2^{-n+1}}\right)\gtrsim_{(1)}\ell\cdot\frac{2^{n}}{\sqrt{d}}\geq_{(2)}C^{\prime}\cdot\sqrt{\ell}~{},$
where $(1)$ follows from the assumption that $n=O(\log d)$, and $(2)$ follows for any constant $C^{\prime}$ of our choice provided that $\ell=C^{\prime\prime}\frac{d}{2^{2n}}$ for a large enough $C^{\prime\prime}$. For sufficiently large $C^{\prime}$, Hoeffding’s bound further ensures that
$\Pr[|\hat{S}^{\ell}\cap S|>3t]>0.999~{},$
@@ Appendix E Proofs for Section 5
We describe an algorithm $\mathcal{A}^{*}$ here: for every cluster $u$ that has been seen at least once in $S$, the algorithm trains a binary-classification model to distinguish $\mathcal{P}_{\theta_{u}}$ from $\mathcal{P}_{0}$ with error at most $\frac{1}{k^{2}}$ (this is possible thanks to Assumption 5.2). Then, upon receiving a query $X$, the algorithm tests whether $X$ was likely from $\mathcal{P}_{\theta_{u}}$ for every seen cluster $u$. $\mathcal{A}^{*}$ outputs $1$ if at least one of the tests outputs $1$. and 00 otherwise. By simple union bound, for $X\sim\mathcal{P}_{0}$, the algorithm makes an error on $X$ with probability at most $O(\frac{1}{k})$. For $X\sim\mathcal{P}_{\theta_{u}}$ where $u$-th cluster is present in $S$, the probability that $\mathcal{A}^{*}$ misclassifies $X$ is at most $\frac{1}{k^{2}}$. For every $X\sim\mathcal{P}_{\theta_{u}}$ where $u\in Z_{n\#0}$, $\mathcal{A}^{*}$ might fail miserably on $X$. However, by Lemma 5.1, the weight of all such clusters, after conditioning on $S$, is at most $\tau_{0}\cdot|Z_{n\#0}|$. This shows that the error of $\mathcal{A}^{*}$ is upper bounded by the right-hand side of Eq. (12), as claimed.

### v2

@@ D.3 Sparse Boolean memorization upper bound
$\mathrm{mem}_{n}(\mathcal{A},\mathbf{P}_{\mathrm{sB}})=I(h;X_{1:n}\mid\theta)\leq I(h;X_{1:n})\leq H(h)=\widetilde{O}\left(\frac{d}{2^{2n}}\right)~.$
Let $\hat{S}\subset[d]$ be the subset that includes coordinates $j\in[d]$ in which the dataset $X_{1:n}$ is constant, namely for which $(X_{i})_{j}=(X_{i^{\prime}})_{j}$ is the same for all $i,i^{\prime}\in[n]$. Furthermore, let $\hat{S}^{\ell}$ be a uniformly random selection of $\ell$ coordinates in $\hat{S}$, and define
@@
$\ell=\Theta\left(\frac{d}{2^{2n}}\right),~\mbox{and}~~t=\sqrt{2\log(200)\ell}=\Theta(\sqrt{\ell}),$
this classifier has the desired properties. We note that the memorization upper bound readily follows by the description length argument above, since all that needs to be stored are the $\ell$ bits of $(X_{1})_{\hat{S}^{\ell}}$ alongside their corresponding indices in $[d]$, which requires $O(\ell\log(d))$ bits.
It remains to show that $h$ is a high accuracy classifier. We first note that in the null case $\mathcal{P}_{0}=\mathcal{U}(\{\pm 1\}^{d})$ and so it holds that
$\mathop{\mathbb{E}}_{X\sim\mathcal{P}_{0}}\left[\langle X_{\hat{S}^{\ell}},(X_{1})_{\hat{S}^{\ell}}\rangle\right]=0~,$
@@
$\Pr_{X\sim\mathcal{P}_{0}}\left[\langle X_{\hat{S}^{\ell}},(X_{1})_{\hat{S}^{\ell}}\rangle\geq t\right]=\Pr_{X\sim\mathcal{P}_{0}}\left[\langle X_{\hat{S}^{\ell}},(X_{1})_{\hat{S}^{\ell}}\rangle\geq\sqrt{2\log(200)\ell}\right]\leq 0.005~.$ (8)
@@
$\mathop{\mathbb{E}}_{X\sim\mathcal{P}_{\theta}}\left[\langle X_{\hat{S}^{\ell}},(X_{1})_{\hat{S}^{\ell}}\rangle\right]=|\hat{S}^{\ell}\cap S|~.$ (9)
@@
Moreover, by symmetry, choosing $\ell$ coordinates in $\hat{S}$ into $\hat{S}^{\ell}$ is equivalent to sub-sampling $\ell$ coordinates among the constant coordinates, and therefore
$\mathop{\mathbb{E}}\left[|\hat{S}^{\ell}\cap S|\right]=\ell\cdot\Pr[j\in S\mid j\in\hat{S}]=\ell\cdot\left(\frac{\nu}{\nu+(1-\nu)2^{-n+1}}\right)\geq_{(1)}\ell\cdot\frac{2^{n}}{4\sqrt{d}}\geq_{(2)}10\sqrt{2\log(200)}\sqrt{\ell}~,$
where $(1)$ holds as long as $n\leq\log_{2}((1-\nu)\sqrt{d})=O(\log d)$ which we assume, and $(2)$ holds for $\ell=\frac{3200\log(200)d}{2^{2n}}$. Since the random variable $|\hat{S}^{\ell}\cap S|$ as a sum of independent indicators, Hoeffding’s bound further ensures that
$\Pr\left[|\hat{S}^{\ell}\cap S|>3t\right]>0.999~,$
@@ Appendix E Proofs for Section 5
We describe an algorithm $\mathcal{A}^{*}$ here: for every cluster $u$ that has been seen at least once in $S$, the algorithm trains a binary-classification model to distinguish $\mathcal{P}_{\theta_{u}}$ from $\mathcal{P}_{0}$ with error at most $\frac{1}{k^{2}}$ (this is possible thanks to Assumption 5.2). Then, upon receiving a query $X$, the algorithm tests whether $X$ was likely from $\mathcal{P}_{\theta_{u}}$ for every seen cluster $u$. $\mathcal{A}^{*}$ outputs $1$ if at least one of the tests outputs $1$. and 0 otherwise. By simple union bound, for $X\sim\mathcal{P}_{0}$, the algorithm makes an error on $X$ with probability at most $O(\frac{1}{k})$. For $X\sim\mathcal{P}_{\theta_{u}}$ where $u$-th cluster is present in $S$, the probability that $\mathcal{A}^{*}$ misclassifies $X$ is at most $\frac{1}{k^{2}}$. For every $X\sim\mathcal{P}_{\theta_{u}}$ where $u\in Z_{n\#0}$, $\mathcal{A}^{*}$ might fail miserably on $X$. However, by Lemma 5.1, the weight of all such clusters, after conditioning on $S$, is at most $\tau_{0}\cdot|Z_{n\#0}|$. This shows that the error of $\mathcal{A}^{*}$ is upper bounded by the right-hand side of Eq. (10), as claimed.